OpenStudy (anonymous):

Please please please help! Easy question!! Please?!?!? Im desperate!! :( The rocks near the shore between two lighthouses at points A and B make the waters unsafe. The measure of (arc)AXB is 300. Waters inside this arc are unsafe. Suppose you are a navigator on a ship at sea. How can you use the lighthouses to keep the ship in safe waters?

OpenStudy (anonymous):

If it's easy why don't you solve it o.O ;D

OpenStudy (anonymous):

Wouldn't you just have to keep an eye on arc AB since that minor arc is the only safe water? The major arc AXB is 300, so arc AB is 60 since it comes full circle.

OpenStudy (anonymous):

So you're saying to just stay 300 feet away from the lighthouses? @Calcmathlete

OpenStudy (anonymous):

Not necessarily...degrees doesn't measure distance.By the way that you named the arcs, it seems like one side of the lighthouses are safe and the other isn't. Therefore, just stay on the 60º side.

OpenStudy (anonymous):

Do you want a picture? I mean i get it, but do you think it would help you to visualize it better?

OpenStudy (anonymous):

@Calcmathlete

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Its under Task 2 if you dont already know. :)

OpenStudy (anonymous):

Ok. If the ship is out of the circle, I would just go around the circle. If you know the length of the arc in between the lighthouses, you can set a ratio an figure out how far away from the lighthouse you'd have to be. For instance, if the distance in between is 6π meters, then set up the proportion. Arc length/2πr = Angle degree/360 6π/dπ = 60/360 6π/dπ = 1/6 36π = dπ d = 36 In this case, stay away from the lighthouses 36 meters because the diameter is the longest distance within the circle.

OpenStudy (anonymous):

So are your equations really the answer or is it an example?

OpenStudy (anonymous):

An example. Distance between light houses/dπ = Angle degree/360. Find d and that's the distance you should stay away from the lighthouses. I'm pretty sure that's what you'd have to do. If it isn't, I'm sorry, but with so little info in the text book, I can't give you a 100% answer. I'd go with that though.

OpenStudy (anonymous):

Alright. Thank you sooo much for the help!!! Ive been asking this question all day long and havent gotten a responce! :)

OpenStudy (anonymous):

@swiftskier96 how did you get the distance?